Reports 1-1 of 1 Clear search Modify search
kiwamu.izumi - 6:39 Thursday 20 June 2019 (9290) Print this report
SR3 oplev loop not good

Darkhan, Yutaro, Kiwamu,

Careful eyes from an expert on the SR3 oplev loop are needed.

During the commissioning activity tonight, we noticed that the beam spot seen by the AS camera had a large oscillation in yaw at 1-2 Hz. We identified that to be SR3 which had a high peak at 1.4-ish Hz in the yaw motion seen by its optical lever. Reducing the gain for the TM stage only, we confirmed that (1) the peak height reduced and (2) the peak location shifted to a lower frequency. This is an indication that the optical lever loop was actually unstable at around 1.4 Hz causing the gain peaking. For now, we are leaving K1:VIS-SR3_TM_DAMP_Y_GAIN to be 0.2 instead of the nominal value of 1.0.

The attached is a comparison of the yaw spectrum with the TM oplev gain of 0.2 (in pink) and 1.0 (in black).

Images attached to this report
Search Help